# QECO: A QoE-Oriented Computation Offloading Algorithm based on Deep Reinforcement Learning for Mobile Edge Computing

### v2

@@ III-D Problem Formulation
[…]
IV DRL-Based Offloading Algorithm
@@
IV-A DQN-based Approach
We utilize the DQN technique to find the mapping between each state-action pair to Q-values in the formulated MDP. As shown in Fig. 2, each MD $i\in\mathcal{I}$ is equipped with a neural network comprising six layers. These layers include an input layer, an LSTM layer, two dense layers, an advantage-value (A&V) layer, and an output layer. The parameter vector $\theta_{i}$ of MD $i$’s neural network is defined to maintain the connection weights and neuron biases across all layers. For MD $i\in\mathcal{I}$, we utilize the state information as the input of neural network. The state information $\lambda_{i}(t)$, $\delta_{i}^{\text{C}}(t)$, $\delta_{i}^{\text{T}}(t)$, $\phi_{i}(t)$, and $\boldsymbol{\eta}_{i}^{\text{E}}(t-1)$ are directly passed to the dense layer, while the state information $\mathcal{H}(t)$ is first supplied to the LSTM layer and then the resulting output is sent to the dense layer. The role and responsibilities of each layer are detailed as follows.
IV-A1 Predicting Workloads at ENs
@@
IV-A2 State-Action Q-Value Mapping
@@
IV-A3 Dueling-DQN Approach for Q-Value Estimation
@@
$\hskip 26.74565pt\Bigg{(}A_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a};\theta_{i})-{1\over|\mathcal{A}|}\sum\limits_{\boldsymbol{a}^{\prime}\in\mathcal{A}}(A_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}^{\prime};\theta_{i})\Bigg{)},$ (26)
@@
Algorithm 1 QECO Algorithm (Offloading Decision)
1:state space $\mathcal{S}$, action space $\mathcal{A}$
2:MD $i\in\mathcal{I}$ experience
@@
14:              Send $(\boldsymbol{s}_{i}(t^{\prime}),\boldsymbol{a}_{i}(t^{\prime}),\boldsymbol{q}_{i}(t^{\prime}),\boldsymbol{s}_{i}(t^{\prime}+1))$ to EN $j_{i}$;
@@
IV-B QoE-Oriented DRL-Based Algorithm
The QECO algorithm is meticulously designed to optimize the allocation of computational tasks between MDs and ENs. Since the training of neural networks imposes an extensive computational workload on MDs, we enable MDs to utilize ENs for training their neural networks, effectively reducing their computational workload. For each MD $i\in\mathcal{I}$, there is an associated EN, denoted as EN $j_{i}\in\mathcal{J}$, which assists in the training process. This EN possesses the highest transmission capacity among all ENs. We define $\mathcal{I}_{j}\subset\mathcal{I}$ as the set of MDs for which training is executed by EN $j\in\mathcal{J}$, i.e. $\mathcal{I}_{j}=\{i\in\mathcal{I}|j_{i}=j\}$. This approach is feasible due to the minimal information exchange and processing requirements for training compared to MD’s tasks. The algorithms to be executed at MD $i\in\mathcal{I}$ and EN $j\in\mathcal{J}$ are given in Algorithms  1 and  2, respectively. The core concept involves training neural networks with MD experiences (i.e., state, action, QoE, next state) to map Q-values to each state-action pair. This mapping allows MD to identify the action in the observed state with the highest Q-value and maximize its long-term QoE.
IV-B1 Offloading Decision Algorithm at MD $i\in\mathcal{I}$
@@
$\boldsymbol{a}_{i}(t)=\begin{cases}\text{arg $\max_{\boldsymbol{a}\in\mathcal{A}}Q_{i}^{\text{E}}(\boldsymbol{s}_{i}(t),\boldsymbol{a};\theta^{\text{E}}_{i})$,}&\text{w.p. $1-\boldsymbol{\epsilon}$,}\\ \text{pick a random action from $\mathcal{A}$,}&\text{w.p. $\boldsymbol{\epsilon}$,}\end{cases}$ (27)
@@
In the next time slot $t+1$, MD $i$ observes the state $\mathcal{S}_{i}(t+1)$. However, due to the potential for tasks to extend across multiple time slots, QoE $\boldsymbol{q}_{i}(t)$ associated with task $z_{i}(t)$ may not be observable in time slot $t+1$. On the other hand, MD $i$ may observe a group of QoEs associated with some tasks $z_{i}(t^{\prime})$ in time slots $t^{\prime}\leq t$. For each MD $i$, we define the set $\mathcal{F}_{i}^{t}\subset\mathcal{T}$ to denote the time slots during which each arriving task $z_{i}(t^{\prime})$ is either processed or dropped in time slot $t$, as given by:
$\mathcal{F}_{i}^{t}=\bigg{\{}t^{\prime}\bigg{|}\;t^{\prime}\leq t,\;\lambda_{i}(t^{\prime})>0,\;(1-x_{i}(t^{\prime}))\;l_{i}^{\text{C}}(t^{\prime})\;+$
$\hskip 56.9055ptx_{i}(t^{\prime})\sum\limits_{\mathcal{J}}\sum\limits_{n=t^{\prime}}^{t}\mathbbm{1}(z_{i,j}^{\text{E}}(n)=z_{i}(t^{\prime}))\;l_{i,j}^{\text{E}}(n)=t\bigg{\}}.$
@@
Algorithm 2 QECO Algorithm (Training Process)
1:Initialize replay buffer $\mathcal{M}_{i}$ for each MD $i\in\mathcal{I}_{j}$;
2:Initialize $\textit{Net}_{i}^{\text{E}}$ and $\textit{Net}_{i}^{\text{T}}$ with random parameters $\theta_{i}^{\text{E}}$ and $\theta_{i}^{\text{T}}$ respectively, for each MD $i\in\mathcal{I}_{j}$;
3:Set Count := 0
4:while True do $\triangleright$ infinite loop
5:     if receive an UpdateRequest from MD $i\in\mathcal{I}_{j}$ then
6:         Send $\theta_{i}^{\text{E}}$ to MD $i\in\mathcal{I}$;
7:     end if
8:     if an experience $(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t),\boldsymbol{q}_{i}(t),\boldsymbol{s}_{i}(t+1))$ is received
9:     from MD $i\in\mathcal{I}_{j}$ then
10:         Store $(\boldsymbol{s}_{i}(t^{\prime}),\boldsymbol{a}_{i}(t^{\prime}),\boldsymbol{q}_{i}(t^{\prime}),\boldsymbol{s}_{i}(t^{\prime}+1))$ in $\mathcal{M}_{i}$;
11:         Get a collection of experiences $\mathcal{I}$ from $\mathcal{M}_{i}$;
12:         for each experience $i\in\mathcal{I}$ do
13:              Get experience $(\boldsymbol{s}_{i}(n),\boldsymbol{a}_{i}(n),\boldsymbol{q}_{i}(n),\boldsymbol{s}_{i}(n+1))$;
14:              Generate $\hat{Q}_{i,n}^{\text{T}}$ according to (28);
15:         end for
16:         Set vector $\hat{\mathbf{Q}}_{i}^{\text{T}}:=(\hat{Q}^{\text{T}}_{i,n})_{n\in\mathcal{N}}$;
17:         Update $\theta_{i}^{\text{E}}$ to minimize $L(\theta_{i}^{\text{E}}$, $\hat{\mathbf{Q}}_{i}^{\text{T}})$ in (30);
18:         Count := Count + 1;
19:         if mod(Count, ReplaceThreshold) = 0 then
20:              $\theta_{i}^{\text{T}}$ := $\theta_{i}^{\text{E}}$;
21:         end if
22:     end if
23:end while
IV-B2 Training Process Algorithm at EN $j\in\mathcal{J}$
@@
In particular, regarding experience $n$, the target-Q value $\hat{Q}_{i,n}^{\text{T}}$ represents the long-term QoE for action $\boldsymbol{a}_{i}(n)$ under state $\boldsymbol{s}_{i}(n)$. This value corresponds to the QoE observed in experience $n$, as well as the approximate expected upcoming QoE. Based on the set $\mathcal{N}$, the EN trains the MD’s neural network using previous sample experiences. Simultaneously, it updates $\theta^{\text{E}}_{i}$ in $\textit{Net}_{i}^{\text{E}}$ and computes vector $\hat{\mathbf{Q}}_{i}^{\text{T}}=(\hat{Q}^{\text{T}}_{i,n})_{n\in\mathcal{N}}$. The key idea of updating $\textit{Net}_{i}^{\text{E}}$ is to minimize the disparity in Q-values between $\textit{Net}_{i}^{\text{E}}$ and $\textit{Net}_{i}^{\text{T}}$, as indicated by the following loss function:
@@
$L(\theta_{i}^{\text{E}},\hat{\mathbf{Q}}_{i}^{\text{T}})={1\over|\mathcal{N}|}\sum\limits_{n\in\mathcal{N}}\bigg{(}Q_{i}^{\text{E}}(\boldsymbol{s}_{i}(n),\boldsymbol{a}_{i}(n);\theta_{i}^{\text{E}})-\hat{Q}^{\text{T}}_{i,n}\bigg{)}^{2}.$ (30)
@@
In every ReplaceThreshold iterations, the update of $\textit{Net}_{i}^{\text{T}}$ will involve duplicating the parameters from $\textit{Net}_{i}^{\text{E}}$ ($\theta_{i}^{T}=\theta_{i}^{E}$). The objective is to consistently update the network parameter $\theta_{i}^{T}$ in $\textit{Net}_{i}^{\text{T}}$, which enhances the approximation of the long-term QoE when computing the target Q-values in (28).
IV-B3 Computational Complexity
@@
The computational complexity of the QECO algorithm is determined by the number of experiences required to discover the optimal offloading policy. Each experience involves backpropagation for training, which has a computational complexity of $\mathcal{O}(C)$, where $C$ represents the number of multiplication operations in the neural network. During each training round triggered by the arrival of a new task, a sample collection of experiences of size $|\mathcal{N}|$ is utilized from the replay buffer. Since the training process encompasses $N^{\text{ep}}$ episodes and there are $K$ expected tasks in each episode, the computational complexity of the proposed algorithm is $\mathcal{O}(N^{\text{ep}}K|\mathcal{N}|C$), which is polynomial. Given the integration of neural networks for function approximation, the convergence guarantee of the DRL algorithm remains an open problem. In this work, we will empirically evaluate the convergence of the proposed algorithm in Section V-B.
V Performance Evaluation
@@
In this section, we first present the simulation setup and training configuration. We then illustrate the convergence of the proposed DRL-based QECO algorithm and evaluate its performance in comparison to three baseline schemes in addition to the existing work [23].
V-A Simulation Setup

### v3

@@ IV-D Problem Formulation
where $\gamma\in(0,1]$ is a discount factor and determines the balance between instant QoE and long-term QoE. As $\gamma$ approaches 0, the MD prioritizes QoE within the current time slot exclusively. Conversely, as $\gamma$ approaches 1, the MD increasingly factors in the cumulative long-term QoE. The expectation $\mathbb{E}[\cdot]$ is taken into consideration of the time-varying system environments. Solving the optimization problem in (24) is particularly challenging due to the dynamic nature of the network. To address this challenge, we introduce a DRL-based offloading algorithm to learn the mapping between each state-action pair and their long-term QoE.
V DRL-Based Offloading Algorithm
@@
V-A DQN-based Approach
We utilize the DQN technique to find the mapping between each state-action pair to Q-values in the formulated MDP. As shown in Fig. 2, each MD $i\in\mathcal{I}$ is equipped with a neural network comprising six layers. These layers include an input layer, an LSTM layer, two dense layers, an advantage-value (A&V) layer, and an output layer. The parameter vector $\theta_{i}$ of MD $i$’s neural network is defined to maintain the connection weights and neuron biases across all layers. For MD $i\in\mathcal{I}$, we utilize the state information as the input of neural network. The state information $\lambda_{i}(t)$, $\delta_{i}^{\text{C}}(t)$, $\delta_{i}^{\text{T}}(t)$, $\phi_{i}(t)$, and $\boldsymbol{\eta}_{i}^{\text{E}}(t-1)$ are directly passed to the dense layer, while the state information $\mathcal{H}(t)$ is first supplied to the LSTM layer and then the resulting output is sent to the dense layer. The roles and responsibilities of each layer are detailed as follows.
@@ V-A DQN-based Approach
V-A1 Predicting Workloads at ENs
@@
V-A2 State-Action Q-Value Mapping
@@
V-A3 Dueling-DQN Approach for Q-Value Estimation
@@
$\hskip 21.05519pt\Bigg(A_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a};\theta_{i})-{1\over|\mathcal{A}|}\sum\limits_{\boldsymbol{a}^{\prime}\in\mathcal{A}}(A_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}^{\prime};\theta_{i})\Bigg),$ (25)
@@
V-B QoE-Oriented DRL-Based Algorithm
The QECO algorithm is meticulously designed to optimize the allocation of computational tasks between MDs and ENs. Since the training of neural networks imposes an extensive computational workload on MDs, we enable MDs to utilize ENs for training their neural networks, effectively reducing their computational workload. For each MD $i\in\mathcal{I}$, there is an associated EN, denoted as EN $j_{i}\in\mathcal{J}$, which assists in the training process. This EN possesses the highest transmission capacity among all ENs. We define $\mathcal{I}_{j}\subset\mathcal{I}$ as the set of MDs for which training is executed by EN $j\in\mathcal{J}$, i.e. $\mathcal{I}_{j}=\{i\in\mathcal{I}|j_{i}=j\}$. This approach is feasible due to the minimal information exchange and processing requirements for training compared to MD’s tasks. The algorithms to be executed at MD $i\in\mathcal{I}$ and EN $j\in\mathcal{J}$ are given in Algorithms 1 and 2, respectively. The core concept involves training neural networks with MD experiences (i.e., state, action, QoE, next state) to map Q-values to each state-action pair. This mapping allows MD to identify the action in the observed state with the highest Q-value and maximize its long-term QoE.
Algorithm 1 Offloading Decision Algorithm at MD $i\in\mathcal{I}$
1:state space $\mathcal{S}$, action space $\mathcal{A}$, parameters vector $\theta_{i}^{\text{E}}$
2:MD $i\in\mathcal{I}$ experience $(\hskip-0.56905pt\boldsymbol{s}_{i}(\hskip-0.56905ptt\hskip-0.56905pt),\boldsymbol{a}_{i}(\hskip-0.56905ptt\hskip-0.56905pt),\boldsymbol{q}_{i}(\hskip-0.56905ptt\hskip-0.56905pt),\boldsymbol{s}_{i}(\hskip-0.56905ptt\hskip-0.85358pt+\hskip-0.85358pt\hskip-0.56905pt1\hskip-0.56905pt\hskip-0.56905pt)\hskip-0.56905pt)$
@@ V-B QoE-Oriented DRL-Based Algorithm
14:     Send $(\boldsymbol{s}_{i}(t^{\prime}),\boldsymbol{a}_{i}(t^{\prime}),\boldsymbol{q}_{i}(t^{\prime}),\boldsymbol{s}_{i}(t^{\prime}\hskip-2.84526pt+1))$ to EN $j_{i}$;
@@
V-B1 Offloading Decision Algorithm at MD $i\in\mathcal{I}$
@@
[…]
@@
In the next time slot $t+1$, MD $i$ observes the state $\mathcal{S}_{i}(t+1)$. However, due to the potential for tasks to extend across multiple time slots, QoE $\boldsymbol{q}_{i}(t)$ associated with task $z_{i}(t)$ may not be observable in time slot $t+1$. On the other hand, MD $i$ may observe a group of QoEs associated with some tasks $z_{i}(t^{\prime})$ in time slots $t^{\prime}\leq t$. For each MD $i$, we define the set $\mathcal{F}_{i}^{t}\subset\mathcal{T}$ to denote the time slots during which each arriving task $z_{i}(t^{\prime})$ is either processed or dropped in time slot $t$, as given by
$\mathcal{F}_{i}^{t}=\bigg\{t^{\prime}\bigg|\;t^{\prime}\leq t,\;\lambda_{i}(t^{\prime})>0,\;(1-x_{i}(t^{\prime}))\;l_{i}^{\text{C}}(t^{\prime})\;$
$\hskip 48.36967pt+\,x_{i}(t^{\prime})\sum\limits_{j\in\mathcal{J}}\sum\limits_{n=t^{\prime}}^{t}\mathbbm{1}(z_{i,j}^{\text{E}}(n)=z_{i}(t^{\prime}))l_{i,j}^{\text{E}}(n)=t\bigg\}.$
@@
Algorithm 2 Training Process Algorithm at EN $j\in\mathcal{J}$
1:experience $(\hskip-0.56905pt\boldsymbol{s}_{i}(\hskip-0.56905ptt\hskip-0.56905pt),\boldsymbol{a}_{i}(\hskip-0.56905ptt\hskip-0.56905pt),\boldsymbol{q}_{i}(\hskip-0.56905ptt\hskip-0.56905pt),\boldsymbol{s}_{i}(\hskip-0.56905ptt\hskip-0.85358pt+\hskip-0.85358pt\hskip-0.56905pt1\hskip-0.56905pt\hskip-0.56905pt)\hskip-0.56905pt)$ from MD $i\in\mathcal{I}$
2:parameters vector $\theta_{i}^{\text{E}}$
3:Initialize replay buffer $\mathcal{M}_{i}$ for each MD $i\in\mathcal{I}_{j}$;
4:Initialize $\textit{Net}_{i}^{\text{E}}$ and $\textit{Net}_{i}^{\text{T}}$ with random parameters $\theta_{i}^{\text{E}}$ and $\theta_{i}^{\text{T}}$ respectively, for each MD $i\in\mathcal{I}_{j}$;
5:Set Count := 0
6:while True do $\triangleright$ infinite loop
7:  if receive an UpdateRequest from MD $i\in\mathcal{I}_{j}$ then
8:   Send $\theta_{i}^{\text{E}}$ to MD $i\in\mathcal{I}$;
9:  end if
10:  if an experience $(\hskip-0.56905pt\boldsymbol{s}_{i}(\hskip-0.56905ptt\hskip-0.56905pt),\boldsymbol{a}_{i}(\hskip-0.56905ptt\hskip-0.56905pt),\boldsymbol{q}_{i}(\hskip-0.56905ptt\hskip-0.56905pt),\boldsymbol{s}_{i}(\hskip-0.56905ptt\hskip-0.85358pt+\hskip-0.85358pt\hskip-0.56905pt1\hskip-0.56905pt\hskip-0.56905pt)\hskip-0.56905pt)$ is received
11:     from MD $i\in\mathcal{I}_{j}$ then
12:   Store $(\boldsymbol{s}_{i}(t^{\prime}),\boldsymbol{a}_{i}(t^{\prime}),\boldsymbol{q}_{i}(t^{\prime}),\boldsymbol{s}_{i}(t^{\prime}\hskip-2.84526pt+1))$ in $\mathcal{M}_{i}$;
13:   Get a collection of experiences $\mathcal{I}$ from $\mathcal{M}_{i}$;
14:   for each experience $i\in\mathcal{I}$ do
15:     Get experience $(\boldsymbol{s}_{i}(\hskip-0.56905ptn\hskip-0.56905pt),\boldsymbol{a}_{i}(\hskip-0.56905ptn\hskip-0.56905pt),\boldsymbol{q}_{i}(\hskip-0.56905ptn\hskip-0.56905pt),\boldsymbol{s}_{i}(\hskip-0.56905ptn\hskip-1.42262pt+1\hskip-0.56905pt))$;
16:     Generate $\hat{Q}_{i,n}^{\text{T}}$ according to (V-B1);
17:   end for
18:   Set vector $\hat{\mathbf{Q}}_{i}^{\text{T}}:=(\hat{Q}^{\text{T}}_{i,n})_{n\in\mathcal{N}}$;
19:   Update $\theta_{i}^{\text{E}}$ to minimize $L(\theta_{i}^{\text{E}}$, $\hat{\mathbf{Q}}_{i}^{\text{T}})$ in (28);
20:   Count := Count + 1;
21:   if mod(Count, ReplaceThreshold) = 0 then
22:     $\theta_{i}^{\text{T}}$ := $\theta_{i}^{\text{E}}$;
23:   end if
24:  end if
25:end while
V-B2 Training Process Algorithm at EN $j\in\mathcal{J}$
@@
In particular, regarding experience $n$, the target-Q value $\hat{Q}_{i,n}^{\text{T}}$ represents the long-term QoE for action $\boldsymbol{a}_{i}(n)$ under state $\boldsymbol{s}_{i}(n)$. This value corresponds to the QoE observed in experience $n$, as well as the approximate expected upcoming QoE. Based on the set $\mathcal{N}$, the EN computes the vector $\hat{\mathbf{Q}}_{i}^{\text{T}}=(\hat{Q}^{\text{T}}_{i,n})_{n\in\mathcal{N}}$ and trains the MD’s neural network (Lines 11-21 of Algorithm 2) to update the parameter vector $\theta^{\text{E}}_{i}$ in $\textit{Net}_{i}^{\text{E}}$ for the next MD’s UpdateRequest. The key idea of updating $\textit{Net}_{i}^{\text{E}}$ is to minimize the disparity in Q-values between $\textit{Net}_{i}^{\text{E}}$ and $\textit{Net}_{i}^{\text{T}}$, as indicated by the following loss function:
@@
$L\Big(\theta_{i}^{\text{E}},\hat{\mathbf{Q}}_{i}^{\text{T}}\Big)={1\over|\mathcal{N}|}\sum\limits_{n\in\mathcal{N}}\bigg(Q_{i}^{\text{E}}(\boldsymbol{s}_{i}(n),\boldsymbol{a}_{i}(n);\theta_{i}^{\text{E}})-\hat{Q}^{\text{T}}_{i,n}\bigg)^{2}.$ (29)
@@
In every ReplaceThreshold iterations, the update of $\textit{Net}_{i}^{\text{T}}$ will involve duplicating the parameters from $\textit{Net}_{i}^{\text{E}}$ ($\theta_{i}^{T}=\theta_{i}^{E}$). The objective is to consistently update the network parameter $\theta_{i}^{T}$ in $\textit{Net}_{i}^{\text{T}}$, which enhances the approximation of the long-term QoE when computing the target Q-values in (V-B1).
V-B3 Computational Complexity
@@
The computational complexity of the QECO algorithm is determined by the number of experiences required to discover the optimal offloading policy. Each experience involves backpropagation for training, which has a computational complexity of $\mathcal{O}(C)$, where $C$ represents the number of multiplication operations in the neural network. During each training round triggered by the arrival of a new task, a sample collection of experiences of size $|\mathcal{N}|$ is utilized from the replay buffer. Since the training process encompasses $N^{\text{ep}}$ episodes and there are $K$ expected tasks in each episode, the computational complexity of the proposed algorithm is $\mathcal{O}(N^{\text{ep}}K|\mathcal{N}|C$), which is polynomial. Given the integration of neural networks for function approximation, the convergence guarantee of the DRL algorithm remains an open problem. In this work, we will empirically evaluate the convergence of the proposed algorithm in Section VI-B.
VI Performance Evaluation
@@
In this section, we first present the simulation setup and training configuration. We then evaluate the performance of QECO in comparison to three baseline schemes in addition to the existing state-of-the-art works [11], [12]. We further analyze the convergence of our proposed algorithm.
VI-A Simulation Setup